# Scalable Dynamic Embedding Size Search for Streaming Recommendation

### v3

@@ 1. Introduction
Many researchers have delved into mixed-dimension embedding size strategies aimed at minimizing memory usage while preserving the performance of recommender systems (Qu et al., 2023; Lyu et al., 2022; Qu et al., 2022; Liang et al., 2023; Chen et al., 2021; Qu et al., 2024; Liu et al., 2021; Joglekar et al., 2020). These strategies involve assigning different embedding sizes based on the characteristics of users and items, such as their interaction counts. However, these methods are tailored for static and non-streaming settings and the embedding sizes remain fixed after allotment, which largely limits their practical utility in streaming settings for two reasons. Firstly, in streaming environments, where users and items exhibit high dynamism, the embedding sizes need to be continually adjusted to accommodate the evolving user and item behavior. For example, the embedding size of a newly released movie may need to be increased as its frequency rises over time. Maintaining consistent embedding sizes in such a situation will lead to performance degradation because it fails to adopt to the changing frequency levels. Secondly, these methods cannot to assign embedding sizes to newly joined users or items. As a result, implementing these methods in streaming scenarios necessitates retraining them from scratch at the onset of each time segment, introducing significant computational overheads.
@@ 2.3. Embedding Compression in Recommendation
A variety of research endeavors have investigated methods for learning compressed embeddings, serving as a memory-efficient alternative to conventional embedding tables in recommender systems, as outlined in (Zheng et al., 2022; Yin et al., 2024; Tran et al., 2024). Approaches like feature selection search, as exemplified by (Wang et al., 2022; Luo et al., 2019; Lin et al., 2022), focus on training models to learn importance scores for effective filtering of unimportant feature fields. Embedding dimension search methods, illustrated by (Liu et al., 2021; Qu et al., 2022; Kong et al., 2023; Qu et al., 2023, 2024), enhance traditional recommender systems by introducing mixed dimensions for features, departing from uniform sizes. Compositional embedding methods (Liang et al., 2024, 2023) represents users and items with a smaller number of embedding vectors (aka. meta-embeddings) to shrink the embedding size. Elastic embedding (Chen et al., 2021; Zheng et al., 2024) enable users and items to share embedding blocks. The knowledge distillation method (Xia et al., 2022, 2023) compresses a model by having a smaller student model learn from a larger teacher model. SCALL maps each user and item to their optimal embedding size so it belong to category of embedding dimension search methods.
@@ 5.1. Baseline Methods
  • CIESS (Qu et al., 2023): An TD3-based embedding size search algorithm that chooses the embedding sizes from a continuous domain.
@@ 5.2. Datasets and Evaluation Protocols
The proposed and baseline method undergo evaluation on two real-world datasets: Amazon-Book (Ni et al., 2019) and Yelp (Wang et al., 2019), both of which contain user ratings. We apply 30-core and 10-core processing on the two datasets, respectively. All the remaining interactions are treated as positive and are sorted in chronological order. After processing, the former dataset comprises 5,076,971 interactions involving 66,351 users and 57,270 items, while the latter involves 2,762,098 interactions between 99,011 users and 56,441 items. A new data segment is created whenever there are 15,000 users or items are introduced (i.e., $m=15,000$). We partition each data segment into 80% for training and 20% for testing, ensuring that the training data precedes the test data chronologically. As a result, 10 data segments are created for the Yelp dataset and 8 data segments are created for the Amazon-Book dataset.
@@ 6. Conclusion and Future Research
Recommender Systems face a significant challenge in the form of the memory bottleneck resulting from uniform embedding sizes for all users and items. Existing dynamic embedding size strategies aim to minimize memory footprint but do not have any fixed constraint on the parameter budget. To address these issues, we proposed SCALL, an automated policy that dynamically adjusts embedding sizes for each user and item and ensures adherence to a predefined budget. The experimental results on the Yelp and Amazon-Book datasets verified the advantageous efficacy of SCALL. However, performing embedding size assignment based on the user and item frequency is biased against the users and items with low frequency. Moving forward, we will concentrate on fair embedding size allocation that overcomes such frequency bias.

### v4

@@ 1. Introduction
Many researchers have delved into mixed-dimension embedding size strategies aimed at minimizing memory usage while preserving the performance of recommender systems (Qu et al., 2023; Lyu et al., 2022; Qu et al., 2022b; Liang et al., 2023; Chen et al., 2021; Qu et al., 2024; Liu et al., 2021; Joglekar et al., 2020). These strategies involve assigning different embedding sizes based on the characteristics of users and items, such as their interaction counts. However, these methods are tailored for static and non-streaming settings and the embedding sizes remain fixed after allotment, which largely limits their practical utility in streaming settings for two reasons. Firstly, in streaming environments, where users and items exhibit high dynamism, the embedding sizes need to be continually adjusted to accommodate the evolving user and item behavior. For example, the embedding size of a newly released movie may need to be increased as its frequency rises over time. Maintaining consistent embedding sizes in such a situation will lead to performance degradation because it fails to adopt to the changing frequency levels. Secondly, these methods cannot to assign embedding sizes to newly joined users or items. As a result, implementing these methods in streaming scenarios necessitates retraining them from scratch at the onset of each time segment, introducing significant computational overheads.
@@ 2.3. Embedding Compression in Recommendation
A variety of research endeavors have investigated methods for learning compressed embeddings, serving as a memory-efficient alternative to conventional embedding tables in recommender systems, as outlined in (Zheng et al., 2022; Yin et al., 2024; Tran et al., 2024). Approaches like feature selection search, as exemplified by (Wang et al., 2022; Luo et al., 2019; Lin et al., 2022), focus on training models to learn importance scores for effective filtering of unimportant feature fields. Embedding dimension search methods, illustrated by (Liu et al., 2021; Qu et al., 2022b; Kong et al., 2023; Qu et al., 2023, 2024), enhance traditional recommender systems by introducing mixed dimensions for features, departing from uniform sizes. Compositional embedding methods (Liang et al., 2024, 2023) represents users and items with a smaller number of embedding vectors (aka. meta-embeddings) to shrink the embedding size. Elastic embedding (Chen et al., 2021; Zheng et al., 2024) enable users and items to share embedding blocks. The knowledge distillation method (Xia et al., 2022, 2023) compresses a model by having a smaller student model learn from a larger teacher model. SCALL maps each user and item to their optimal embedding size so it belong to category of embedding dimension search methods.
@@ 5.1. Baseline Methods
  • CIESS (Qu et al., 2023): A TD3-based embedding size search algorithm that chooses the embedding sizes from a continuous domain.
@@ 5.2. Datasets and Evaluation Protocols
The proposed and baseline method undergo evaluation on two real-world datasets: Amazon-Book (Ni et al., 2019) and Yelp  [^1] https://business.yelp.com/data/resources/open-dataset/, both of which contain user ratings. We apply 30-core and 10-core processing on the two datasets, respectively. All the remaining interactions are treated as positive and are sorted in chronological order. After processing, the former dataset comprises 5,076,971 interactions involving 66,351 users and 57,270 items, while the latter involves 2,762,098 interactions between 99,011 users and 56,441 items. A new data segment is created whenever 15,000 users or items are introduced (i.e., $m=15,000$). We partition each data segment into 80% for training and 20% for testing, ensuring that the training data precedes the test data chronologically. As a result, 10 data segments are created for the Yelp dataset and 8 data segments are created for the Amazon-Book dataset.
@@ 6. Conclusion and Future Research
Recommender Systems face a significant challenge in the form of the memory bottleneck resulting from uniform embedding sizes for all users and items. Existing dynamic embedding size strategies aim to minimize memory footprint but do not have any fixed constraint on the parameter budget. To address these issues, we proposed SCALL, an automated policy that dynamically adjusts embedding sizes for each user and item and ensures adherence to a predefined budget. The experimental results on the Yelp and Amazon-Book datasets verified the advantageous efficacy of SCALL. However, performing embedding size assignment based on the user and item frequency is biased against the users and items with low frequency. Moving forward, we will concentrate on fair embedding size allocation that overcomes such frequency bias using other importance signals such as model confidence (Qu et al., 2021, 2022a).